# Analog Spiking Neuron in CMOS 28 nm Towards Large-Scale Neuromorphic Processors

### v1

@@ II-D ASIC implementation
Figure 2: LIF neuron layout for implementation in TSMC 28 nm CMOS.
The proposed LIF neuron design was fabricated using the TSMC 28 nm CMOS technology node. Fig. 2 illustrates the circuit layout, which occupies an area of approximately $5~{}\mu m$ by $6~{}\mu m$. The total chip implementation has a total area of $1000~{}\mu m$ by $1000~{}\mu m$ and is encapsulated in a dual inline 20 package.
To facilitate testing after fabrication, all the inputs, outputs, and power connections are linked to the ASIC’s pads with Electro-Static Discharge (ESD) protection.
@@ III-B Software Emulation
Our evaluation methodology also uses Post-Training Quantization (PTQ). The model is initially trained with half-precision floating-point weights and subsequently quantized to 4-bit resolution. We then incorporate the experimentally-derived activation function from chip measurements for accuracy evaluation. This approach simultaneously enhances computational efficiency and emulates the model’s performance within a simulated resource-constrained environment. Such emulation is crucial for assessing the architecture’s real-time applicability and scalability for edge AI applications
$V_{\text{reset}},V_{\text{th}},V_{\text{spike}},\tau_{m},R_{m},dt,t_{\text{refractory}},N_{\text{syn}},W_{i,j},I_{\text{ref}},N_{i}$ (number of neurons in layer $l_{i}$)
for each layer $l_{i}$ do
     for each neuron $k$ in layer $l$ do
         $V_{\text{mem}}[k,l]\leftarrow V_{\text{reset}}$
         $t_{\text{last\_spike}}[k,l]\leftarrow-\infty$triangleright Initialize last spike time
         for each time step $t$ do
              $I_{\text{synaptic}}[k,l](t)\leftarrow\sum_{i}^{N_{i}}W_{i,j}[k]\times I_{\text{ref}}$
              if $(t-t_{\text{last\_spike}}[k,l])\geq t_{\text{refractory}}$ then
                  $V_{\text{mem}}(t+1)\leftarrow V_{\text{mem}}(t)+dt\times\left(-\frac{(V_{\text{mem}}(t)-V_{\text{reset}})+R_{m}\times I_{\text{synaptic}}[k,l](t)}{\tau_{m}}\right)$
                  if $V_{\text{mem}}(t+1)\geq V_{\text{th}}$ then
                       $V_{\text{mem}}[k,l]\leftarrow V_{\text{reset}}$
                       $V_{\text{spike}}[k,l]=1$
                       $t_{\text{last\_spike}}[k,l]\leftarrow t$
                  else
                       $V_{\text{spike}}[k,l]=0$
                  end if
              end if
         end for
     end for
end for
Algorithm 2 Fully-Connected SNN Computation Flow using LIF Model
@@
The surrogate gradient approach addresses the problem by keeping the Heaviside function during the forward pass, and computing its derivative with a continuous approximation during the backward pass. the $Softmax$ function can be used for the purpose, with a parameter controlling its steepness near the Heaviside threshold. [30].
@@ IV Results
We present here the measurement results of the implemented LIF neuron’s characteristics with regard to the integration shape, the spiking behavior and frequency, the standby power, and the energy dissipation per spike.
@@ IV-A LIF Synaptic Current Generation
The primary challenges in characterizing this architecture involve accurately measuring input current $I_{IN}$ at picoampere levels. To achieve this, the methodology proposed in this paper employs a combination of measurement and post-layout simulation to determine the input current. The input current to the LIF is sourced from a voltage source provided by the picoammeter connected to node $V_{IN}$ in Fig. 8. The voltage is applied to the shorted drain and gate of transistor $M_{1}$ via a 200 ohm resistor ($R_{ESD}$) in the ESD protection circuit. In the next paragraph, $I_{IN-sim}$ is the simulated $I_{IN}$ and $I_{IN-exp}$ is the measured $I_{IN}$ using the experimental setup.
@@
Figure 9: Quantitative analysis of the deviation between simulated and measured curves highlights the accuracy of the co-simulation measurement approach, further confirming its validity.
@@ IV-B Membrane Voltage Behavior over Time
The exponential rise in current increases the charging rate of the gate capacitance of transistor $M_{3}$. This variable charging rate results in different activation timings for $M_{3}$, which in turn rapidly up-swings the membrane voltage $V_{mem}$. Consequently, $V_{mem}$ surpasses the inverter threshold $V^{M5-M6}_{INV}$, triggering a spike at the output. Following this spike, the rest of the circuit, particularly transistor $M_{4}$, resets $V{mem}$ back to $V_{reset}$.
@@ IV-C Spiking Behavior over Time
The figure also shows a linear increase of the spike peak with the input current level, although not projected in equation 13. The reason is that given the technology node and the small transistor ratio’s chosen, the speed of the reset block dominates the recursive link between the spike generation inverters and the reset transistors.As varying the input current level leads to different integration slopes in capacitor $C_{mem}$, different delays occur before $V_{mem}$ reaches the switching threshold of inverter pair $M_{5}$-$M_{6}$. Then, $C_{mem}$ keeps charging and, after inverter pair $M_{7}$-$M_{8}$ also switches State, $C_{res}$ starts charging as well until it voltage reaches the threshold level to reset $C_{mem}$ through transistor pair $M_{5}$-$M_{6}$. Given the faster rise of $V_{mem}$ for larger input currents, a higher value of $V_{OH}$ across transistor $M_{8}$ is reached before $C_{mem}$ reset, hence the observed variation of the peak voltage with the intensity of the input current.
@@ IV-D Spiking Frequency
Fig. 12 shows the firing frequency of the implemented LIF neuron for different input synaptic current amplitudes. The neuron’s firing frequency increases with the input synaptic current, going from 10 kHz to 350 kHz in the 10 pA to 3000 pA input range, respectively. These results were collected from 20 ASIC samples randomly selected that were wire-bonded to a Dual-In-Line 20 (DIP20) package in order to test them using the PCB test bench.
Figure 12: Measured spiking frequency for input synaptic current levels from 0 pA to 3000 pA, showing the relationship between the spike generation rate and the intensity of synaptic current.
@@ IV-E Energy Consumption per Spike Event
Figure 13: Measured energy consumption per spike as a function of the input synaptic current for 20 different chip samples.
@@ V Discussion
This paper presents an implementation of a spiking neuron in a 28 nm CMOS node. The proposed design has been fabricated and comprehensively characterized. Based on this detailed characterization, a model is developed using PyTorch to enable practical emulation on GPUs at a very high neuron density.
In the proposed LIF architecture, the leak component isn’t explicitly represented by a lumped model or transistors but is instead inherent through the parasitic leakage currents of transistor $M_{4}$.
The comparison in Table II of our LIF design with published spiking neuron implementations in CMOS technology shows the achievement of our objective of a candidate for a large-scale neural network with minimal silicon footprint and energy consumption. This is shown clearly in Fig. 14 which provides a scatter diagram of recent spiking neuron implementations in CMOS technology, comparing their energy consumption per spike with their active area.
@@
The results suggest that the proposed CMOS implementation of the LIF spiking neuron is a promising avenue for neuromorphic computing. A large-scale implemenations can be emulated by successfully deriving bio-mimetic models from analog neurons. This approach provides a robust method for verifying NeuroSOC designs before mass production and highlights the potential for creating efficient and scalable neuromorphic computing systems.
The results of this paper showcase the effective implementation and characterization of the proposed LIF architecture, which accurately integrated input synaptic currents over time, resulting in a spiking behavior that depended on the synaptic current intensity in an increasing fashion. Notably, the energy consumption per spike event was found to depend in a pseudo-linear manner on the input current level. However, there are also inherent limitations of the proposed method. The most significant are the exclusion of energy costs associated with interconnections, synaptic integration, and supporting circuits such as communication protocols. These factors, along with potential delays in weight updates and data transmission challenges, might impact the overall accuracy and efficiency of large-scale neuromorphic systems. Still, the work offers a useful foundation for future research and development in the field, especially in scaling up the LIF neuron design for extensive neuromorphic implementations.
The argument could also be made that the synaptic connections and integration are not considered in the proposed emulation method. Another argument is that the synapses are the bottleneck in a fully connected configuration. On the other hand, memristor technology is developing fast and their active footprint on silicon is much smaller than CMOS neurons. However, memristor-based devices have still a long way to go to achieve the necessary precision, reliability and robustness against PVT that will enable large-scale implementation of NeuroSoCs.
Additionally, the primary focus of this paper is to propose and showcase a scalable LIF architecture designed for CMOS implementation, aimed at facilitating the development of large-scale neuromorphic chips. The integration of synapses is out of the scope of this work.
@@
Furthermore, the energy consumption estimate for a full neural network inference is computed based on the experimental measurement per spike. To do this estimation, the pre-trained neural network is run on the test fraction of the dataset and the number of spikes produced by each neuron in every layer is determined to obtain the total number of generated spikes by the network during one inference. Then, this value is multiplied by the average measured energy per spike of 2 fJ/Spike. The recorded results are presented in Table I. This methodology offers valuable insights into emulating and forecasting the suitability of LIF neuron design for large-scale neuromorphic systems.
There are still parameters that remain to be factored in such as potential delays in weight updates, processing of input data, and the transmission of network outputs off-chip through communication protocols such as I2C or SPI, and challenges like communication errors due to data links and noise from power sources further add to the complexity of the NeuroSoCs. In this respect, comprehensive testing and large-scale trials are necessary to fully grasp how these factors might influence the accuracy and energy efficiency of a large-scale neuromorphic system.

### v2

@@ II-D ASIC implementation
Figure 2: Layout implementation of the LIF neuron in TSMC 28 nm CMOS technology. The figure illustrates only the routing metals, excluding transistor-level details, due to non-disclosure agreements (NDAs) with the foundry. Post-layout simulations estimate gate-leakage currents to range from 100 fA to 1 pA under typical operating conditions, rendering them negligible compared to Isyn and Imem. The design complies with the manufacturer’s design rule checking (DRC) guidelines, including the minimum required lengths for gate polysilicon
The proposed LIF neuron design was fabricated using the TSMC 28 nm CMOS technology node. The choice of the 28 nm TSMC CMOS technology for the silicon neuron design was driven by its unique advantages for neuromorphic circuit implementations. This technology node enables significant reductions in energy consumption and area footprint due to its smaller feature size and lower capacitance. Moreover, in smaller technology nodes such as 28 nm, transistors can be operated at lower core voltages, and logic gates can function effectively at reduced power supply voltages. These factors significantly lower overall power consumption, which is particularly advantageous for large-scale implementations in spiking neural networks. Additionally, the 28 nm process provides access to advanced features such as high-speed and low-leakage transistors, which are crucial for designing efficient and scalable neuromorphic circuits. These characteristics make the 28 nm technology particularly well-suited for the development of silicon neurons that can be integrated into large-scale spiking neural networks with minimal resource constraints
Fig. 2 illustrates the circuit layout, which occupies an area of approximately $5~\mu m$ by $6~\mu m$. The total chip implementation has a total area of $1000~\mu m$ by $1000~\mu m$ and is encapsulated in a dual inline 20 package. To facilitate testing after fabrication, all the inputs, outputs, and power connections are linked to the ASIC’s pads with Electro-Static Discharge (ESD) protection.
@@ III-A2 LIF Spiking Behavior
The very low voltage smooth pulses generated by the silicon neuron circuit are destined to be interfaced with digital circuits using a custom inverter designed with a low triggering threshold. This inverter will translate the pulses from the low-voltage power supply domain to the digital gates domain, ensuring compatibility with standard digital circuitry. The $1\;\mu s$ duration of the spike responses is maintained during this process, allowing precise temporal encoding to be preserved. Once translated, the pulses can be processed by subsequent digital circuitry for integration into larger-scale systems
@@ III-B Software Emulation
Our evaluation methodology also uses Post-Training Quantization (PTQ). The model is initially trained with half-precision floating-point weights and subsequently quantized to 4-bit resolution. We then incorporate the experimentally-derived activation function from chip measurements for accuracy evaluation. This approach simultaneously enhances computational efficiency and emulates the model’s performance within a simulated resource-constrained environment. Such emulation is crucial for assessing the architecture’s real-time applicability and scalability for edge AI applications.
@@
The surrogate gradient approach addresses the problem by keeping the Heaviside function during the forward pass, and computing its derivative with a continuous approximation during the backward pass. the $Softmax$ function can be used for the purpose, with a parameter controlling its steepness near the Heaviside threshold. [35]. The MNIST-trained SNN uses the measured f-I curves of the fabricated neuron as activation functions to closely emulate the circuit’s behavior. Surrogate gradient training was employed to address the non-differentiable nature of spike generation in the temporal domain. While rate-coding networks do not typically require surrogate gradients, this paper use of spike-latency coding for training necessitated this approach to enable efficient temporal learning.
The proposed LIF neuron circuit is intended as a scalable building block for neuromorphic systems leveraging efficient spiking neural network (SNN) architectures. While the MNIST emulation experiment using the the measured f-I curves employs a rate-coding approach as a proof-of-concept, the circuit is designed to support advanced coding schemes, such as spike-latency or event-based coding, which maximize the temporal precision of spiking neurons and minimize energy consumption by reducing the number of spikes required for computation.
@@ IV Results
We present here the measurement results of the implemented LIF neuron’s characteristics with regard to the integration shape, the spiking behavior and frequency, the standby power, and the energy dissipation per spike. Beyond mere footprint reduction, our sub-2 fJ/spike performance arises from a combination of design and technology choices. First, by employing minimum-length channels in TSMC’s 28 nm node and biasing key transistors in deep sub-threshold, we realize picoampere-scale currents and negligible leakage, directly minimizing dynamic energy. Second, the use of an ultra-small membrane capacitance ($C_{mem}=3.47\ fF$) lowers the charge per spike ($Q=C_{mem}V_{DD}$), shrinking both stored and discharged energy. Third, the neuron’s capability to fire at up to 300kHz300\ kHz means each high-power switching event occupies only a brief interval, and since $E_{spike}=P_{avg}/f$, a higher frequency directly drives down energy per event. Finally, operating at a supply voltage of only $V_{DD}=250\ mV$ exploits the quadratic dependence of dynamic power ($P_{dyn}\propto V_{DD}^{2}$) to curtail both static and dynamic dissipation. Together, these factors—sub-threshold transistor operation, minimal $C_{mem}$, high spiking rate, and low $V_{DD}$—enable our analog LIF neuron to achieve energy efficiency alongside its compact $34\ \mu m^{2}$ area.
@@ IV-A LIF Synaptic Current Generation
The primary challenges in characterizing this architecture involve accurately measuring input current $I_{IN}$ at picoampere levels. To achieve this, the methodology proposed in this paper is based on a combination of measurement and post-layout simulation to determine the input current. The input current to the LIF is sourced from a voltage source provided by the picoammeter connected to node $V_{IN}$ in Fig. 8. The voltage is applied to the shorted drain and gate of transistor $M_{1}$ via a 200 ohm resistor ($R_{ESD}$) in the ESD protection circuit. In the next paragraph, $I_{IN-sim}$ is the simulated $I_{IN}$ and $I_{IN-exp}$ is the measured $I_{IN}$ using the experimental setup.
@@
This method characterizes the relationship between the LIF neuron’s firing rate and its input synaptic current. Since no CMOS synapses are involved, the only relevant noise sources during current generation lie along the path from node $V_{IN}$ to node $V_{MEM}$ (Fig. 8). Within this path, the principal contributors are the thermal and flicker noise of transistors M1 and M2, which form the current mirror. As these transistors integrate current onto $C_{MEM}$, their noise is likewise integrated and averaged over many spikes. Leakage currents at the gates, drains, and parasitic capacitances of M2–M6 contribute similarly minor noise. Although we cannot directly measure these picoampere-level noise contributions, post‐layout noise simulations confirm they do not materially affect performance. Suppressing such noise would require specialized processes (e.g., FDSOI) or cryogenic operation, which are beyond this work’s scope. Finally, spiking neural networks are inherently robust to input variability, so even with the inherent noise of the M1–M2 current mirror, the proposed design will scale effectively to large SNNs.
Figure 9: Quantitative analysis of the deviation between simulated and measured curves highlights the accuracy of the co-simulation measurement approach, further confirming its validity. (a) Depicts the spiking frequency of the LIF neuron at node $V_{out}$ in respect of the input voltage from the Keithley pico-ammeter $V_{IN}$. (b) Depicts the input current at node $I_{IN}$ depicted in Fig. 8 in respect of the input voltage from the Keithley pico-ammeter $V_{IN}$. (c) Depicts the post-layout simulation results of the input synaptic current $I_{SYN}$ in respect of the input voltage from the Keithley pico-ammeter $V_{IN}$. (d) Depicts the spiking frequency of the LIF neuron at node $V_{out}$ in respect of the input synaptic current $I_{SYN}$.
@@ IV-B Membrane Voltage Behavior over Time
The exponential rise in current increases the charging rate of the gate capacitance of transistor $M_{3}$. This variable charging rate results in different activation timings for $M_{3}$, which in turn rapidly up-swings the membrane voltage $V_{mem}$. Consequently, $V_{mem}$ surpasses the inverter threshold $V^{M5-M6}_{INV}$, triggering a spike at the output. Following this spike, the rest of the circuit, particularly transistor $M_{4}$, resets VmemV{mem} back to $V_{reset}$.
The selection of the membrane capacitance $C_{mem}$ directly governs the integration rate and hence the neuron’s spiking frequency. From Eq. (3), the integration slope is given by $dV_{mem}/dt=I_{mem}/C_{mem}$, so a smaller $C_{mem}$ produces a steeper voltage ramp, reducing the inter-spike interval and enabling higher maximum firing rates. Conversely, according to the energy relation $E_{spike}\approx\tfrac{1}{2}C_{mem}V_{DD}^{2}$, a larger capacitance increases the energy stored and dissipated per spike. Thus, $C_{mem}$ must be carefully sized to balance speed, energy consumption, and stability: excessively small capacitances yield faster spikes but amplify noise and process-variation sensitivity, while overly large capacitances improve robustness but incur higher energy costs and slower operation. In our design, $C_{mem}=3.47\,fF$ was chosen to support >> 300 kHz firing at $I_{syn}$ up to 3 nA, while keeping $E_{spike}$ below 2 fJ.
@@ IV-C Spiking Behavior over Time
The figure also shows a linear increase of the spike peak with the input current level as anticipated by equation Eq. 16. The reason is that given the technology node and the small transistor ratio’s chosen, the speed of the reset block dominates the recursive link between the spike generation inverters and the reset transistors.As varying the input current level leads to different integration slopes in capacitor $C_{mem}$, different delays occur before $V_{mem}$ reaches the switching threshold of inverter pair $M_{5}$-$M_{6}$. Then, $C_{mem}$ keeps charging and, after inverter pair $M_{7}$-$M_{8}$ also switches State, $C_{res}$ starts charging as well until it voltage reaches the threshold level to reset $C_{mem}$ through transistor pair $M_{5}$-$M_{6}$. Given the faster rise of $V_{mem}$ for larger input currents, a higher value of $V_{OH}$ across transistor $M_{8}$ is reached before $C_{mem}$ reset, hence the observed variation of the peak voltage with the intensity of the input current.
@@ IV-D Spiking Frequency
Fig. 12 shows the firing frequency of the implemented LIF neuron for different input synaptic current amplitudes. The neuron’s firing frequency increases with the input synaptic current, going from 10 kHz to 300 kHz in the 10 pA to 10000 pA input range, respectively. These results were collected from 20 ASIC samples randomly selected that were wire-bonded to a Dual-In-Line 20 (DIP20) package in order to test them using the PCB test bench.
Figure 12: Measured spiking frequency for input synaptic current levels from 0 pA to 10000 pA, showing the relationship between the spike generation rate and the intensity of synaptic current.
@@ IV-E Energy Consumption per Spike Event
Figure 13: Measured energy consumption per spike as a function of the input synaptic current.
@@ V Discussion
This paper presents an implementation of a spiking neuron in a 28 nm CMOS node. The proposed design has been fabricated and comprehensively characterized. Based on this detailed characterization, a model is developed using PyTorch to enable practical emulation on GPUs at a large neuron density. The comparison in Table II of our LIF design with published spiking neuron implementations in CMOS technology shows the achievement of our objective of a candidate for a large-scale neural network with minimal silicon footprint and energy consumption. This is shown clearly in Fig. 14 which provides a scatter diagram of recent spiking neuron implementations in CMOS technology, comparing their energy consumption per spike with their active area.
@@
To the best of the authors knowledge, no prior work has combined operation at a 250 mV supply, a per-spike energy below 2 fJ, and an active area under 35 $\mu m^{2}$. Fabricated in TSMC 28 nm CMOS, the proposed neuron achieves 1.61 fJ/spike in only 34 $\mu m^{2}$ at 250 mV, representing an improvement in energy density over the closest analog designs and a reduction in active area compared to early 65 nm implementations, while still supporting spiking frequencies up to 300 kHz. This combination of scaling, sub-threshold integration, low-voltage operation and compact layout establishes a new performance benchmark for ultra-low-power, high-density analog LIF neurons. The results suggest that the proposed CMOS implementation of the LIF spiking neuron is a promising avenue for neuromorphic computing. A large-scale implemenations can be emulated by successfully deriving bio-mimetic models from analog neurons. This approach provides a robust method for verifying NeuroSOC designs before mass production and highlights the potential for creating efficient and scalable neuromorphic computing systems. The results of this paper showcase the effective implementation and characterization of the proposed LIF architecture, which accurately integrated input synaptic currents over time, resulting in a spiking behavior that depended on the synaptic current intensity in an increasing fashion. Notably, the energy consumption per spike event was found to depend in a pseudo-linear manner on the input current level. However, there are also inherent limitations of the proposed method. It is worth noting that the proposed method does not include the integration of synaptic connections, which are often considered a bottleneck in fully connected configurations. While memristor technology offers advantages such as a smaller active footprint compared to CMOS neurons, it still faces significant challenges in achieving the precision and reliability necessary for large-scale NeuroSoC implementations. The primary focus of this paper is to propose and validate a scalable LIF architecture specifically designed for CMOS implementation. The integration of synapses, while crucial for complete neuromorphic systems, falls outside the scope of this work and will be considered in future extensions.
Fabricating an analog LIF neuron in TSMC’s 28 nm node posed several interrelated challenges. Operating key transistors in deep sub-threshold to achieve picoampere currents demanded careful matching addressed via common-centroid, multi-finger layouts. The 28 nm design rules yielded a layout mask that can inflate parasitic capacitances. To mitigate this, we performed extensive post-layout and corners simulations at FF through SS corners. Guided by these results, we tuned transistor aspect ratios and bias voltages to secure reliable firing behavior up to 300 kHz spiking and under 2 femto-Joules per spike. This methodology ensures a PVT reliable design that achieves its low-voltage, ultra-low-energy, and compact-area targets.
@@
Furthermore, the energy consumption estimate for a full neural network inference is computed based on the experimental measurement per spike. To do this estimation, the pre-trained neural network is run on the test fraction of the dataset and the number of spikes produced by each neuron in every layer is determined to obtain the total number of generated spikes by the network during one inference. Then, this value is multiplied by the average measured energy per spike of 2 fJ/Spike. The recorded results are presented in Table I. This methodology offers valuable insights into emulating and forecasting the suitability of LIF neuron design for large-scale neuromorphic systems. While the energy consumption estimates presented in this work focus on the LIF neuron circuitry, it is important to note that synapses typically contribute significantly more to the total power consumption in neuromorphic systems, as observed in recent works [10, 8] . Our estimates therefore provide an evaluation of the neuron’s energy efficiency in isolation, which was the primary focus of this study. Future work will aim to extend these analyses to include synaptic circuits to provide a more comprehensive understanding of the energy profile of the complete system.
The MNIST-trained neural network was included as a proof-of-concept to demonstrate the applicability of the proposed LIF neuron for practical spiking neural network (SNN) applications. The primary objective was to validate the feasibility of implementing such networks using the characteristics of the proposed silicon neuron. As this paper primarily focuses on hardware design and characterization, a detailed exploration of training metrics, such as learning rate, hyperparameters, accuracy over epochs, and dataset splits, was not the main scope of this work. These aspects will be investigated in greater depth in future studies focused on the deployment and optimization of SNNs for specific tasks. The observed performances figures are not intended to compete with state-of-the-art results but to demonstrate that the circuit can support spike-based computation in real-world tasks. The reduced accuracy is attributed to the simplified network architecture, lack of inhibitory mechanisms, and the use of experimentally measured circuit parameters, which were not fine-tuned for optimal performance
The energy-saving potential of spiking neural networks (SNNs) lies primarily in their ability to exploit efficient coding schemes, such as spike-latency or event-based coding, which minimize the number of spikes required for computation and take advantage of temporal sparsity. While SNNs can operate in a rate-coding context, this approach often negates many of the energy and computational benefits, as it requires temporal averaging to estimate rates and results in a linear relationship between spike count and accuracy. In contrast, binary or sparse coding schemes provide logarithmic scaling of spike count with accuracy, significantly improving efficiency. The MNIST experiment in this work employed a rate-coding approach for simplicity and proof-of-concept validation; however, the proposed neuron circuit is designed to support more efficient coding schemes in future implementations.
At the circuit level, the design departs from prior analog LIF implementations. First, the leak conductance is implemented implicitly via the sub-threshold operation of M3–M4, eliminating the need for a large on-chip resistor and reducing silicon area. Second, the integration block (M1–M2) operates at only 250 mV supply, extending sub-threshold current‐mirror operation into a voltage regime rarely explored by existing designs. Third, the reset mechanism employs capacitive coupling (Cres) and a threshold‐triggered transistor M4 to achieve a precise 1 µs spike width without additional digital control. Fourth, the two‐stage inverter chain (M5–M8) is optimized for sub-threshold switching, using a low inversion threshold (appx. 60 mV) and minimal voltage swing to further cut dynamic energy.